# HOG-Diff: Higher-Order Guided Diffusion for Graph Generation

### v1

@@ 3 Higher-order Guided Diffusion Model
In the following, we provide supportive theoretical evidence for the efficacy of HOG-Diff, demonstrating that the proposed framework achieves faster convergence in score-matching and tighter reconstruction error bounds compared to standard graph diffusion works.
Proposition 3.3 (Informal).
@@
Suppose the loss function $\ell^{(k)}(\bm{\theta})$ in Eq. 11 is $\beta$-smooth and satisfies the $\mu$-PL condition in the ball $B\left(\bm{\theta}_{0},R\right)$. Then, the expected loss at the $i$-th iteration of the training process satisfies:
$\mathbb{E}\left[\ell^{(k)}(\bm{\theta}_{i})\right]\leq\left(1-\frac{b\mu^{2}}{\beta N(\beta N^{2}+\mu(b-1))}\right)^{i}\ell^{(k)}\left(\bm{\theta}_{0}\right),$
@@
Following Luo et al. (2023a), we define the expected reconstruction error at each generation process as $\mathcal{E}(t)=\mathbb{E}\left\|\bar{\bm{G}}_{t}-\widehat{\bm{G}}_{t}\right\|^{2}$, where $\bar{\bm{G}}_{t}$ represents the data reconstructed sing the ground truth score $\nabla\log p_{t}(\cdot)$ and $\widehat{\bm{G}}_{t}$ denotes the data reconstructed with the learned score function $\bm{s}_{\bm{\theta}}$. We establish that the reconstruction error in HOG-Diff is bounded more tightly than in classical graph generation models, ensuring superior sample quality.
Proposition 3.4.
@@
where $\alpha(0)=C^{2}\ell^{(1)}(\bm{\theta})\int_{0}^{\tau_{1}}g_{1,s}^{4}\mathop{}\!\mathrm{d}{s}+C\mathcal{E}(\tau_{1})\int_{0}^{\tau_{1}}h_{1,s}^{2}\mathop{}\!\mathrm{d}{s}$, $\gamma(s)=C^{2}g_{1,s}^{4}\|\bm{s}_{\bm{\theta}}(\cdot,s)\|_{\mathrm{lip}}^{2}+C\|h_{1,s}\|_{\mathrm{lip}}^{2}$, and $h_{1,s}=\theta_{s}\left(1+\frac{2}{e^{2\bar{\theta}_{s:\tau_{1}}}-1}\right)$. Furthermore, we can derive that the reconstruction error bound of HOG-Diff is sharper than classical graph generation models.
The propositions above rely primarily on mild assumptions, such as smoothness and boundedness, without imposing strict conditions like the target distribution being log-concave or satisfying the log-Sobolev inequality. Their formal statements and detailed proofs are postponed to App. A. We experimentally verify these Propositions in Sec. 4.3.
@@ 4 Experiments
We assess HOG-Diff against state-of-the-art baselines for both molecular and generic graph generation. Further ablation studies are conducted to analyze the impact of different topological guides. More details can be found in App. F.
@@ 4.1 Molecule Generation
Experimental Setup
To assess the capability of the proposed method in molecular generation, we conduct evaluations on two well-known molecular datasets: QM9 (Ramakrishnan et al., 2014) and ZINC250k (Irwin et al., 2012), and obtain the intermedia higher-order skeletons using the 2-cell complex filtering. We evaluate the quality of 10, 000 generated molecules with six metrics as in Jo et al. (2022): Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) MMD (Costa & Grave, 2010), Fréchet ChemNet Distance (FCD) (Preuer et al., 2018), Validity (Val.), Validity without correction (Val. w/o corr.), Uniqueness (Uni.), and Novelty (Nov.) (Jo et al., 2022).
Baselines
We evaluate our model against state-of-the-art molecular generation models, including auto-regressive methods, GraphAF (Shi et al., 2020), and GraphDF (Luo et al., 2021). For a fair comparison, as recommended by Jo et al. (2022), we extend GraphAF and GraphDF to account for formal charges in the molecular generation, termed GraphAF+FC and GraphDF+FC, respectively. We also compare our HOG-Diff with various flow-based and diffusion-based methods, including MoFlow (Zang & Wang, 2020), EDP-GNN (Niu et al., 2020), Graph-EBM (Liu et al., 2021), GDSS (Jo et al., 2022), and DiGress (Vignac et al., 2023).
Sampling Quality
We visualize the molecule generation process in Fig. 3 with more examples deferred to App. G. It can be observed that our model explicitly preserves higher-order structures during the generation process. Tab. 1 indicates that HOG-Diff consistently outperforms both auto-regressive and one-shot models. Notably, the dramatic decrease in NSPDK and FCD implies that HOG-Diff is able to generate molecules with data distributions close to those of the real molecules in both the chemical and graph space.
Figure 3: Visualization of molecular graphs at different stages of the reverse generative process. Model trained on Zinc250k.
@@ 4.2 Generic Graph Generation
Table 2: Generation performance on generic graph datasets. The best results are highlighted in bold and an underline indicates the second-best result. Hyphen (-) denotes that the results are not provided in the original paper.
Methods Community-small Enzymes Ego-small
@@
Experimental Setup
To display the topology distribution learning ability, we access HOG-Diff over three common generic graph datasets: (1) Community-small, containing 100 randomly generated community graphs; (2) Ego-small, comprising 200 small ego graphs derived from the Citeseer network dataset; (3) Enzymes, featuring 587 protein graphs representing tertiary structures of enzymes from the BRENDA database. Intermediate higher-order skeletons are obtained through 3-simplicial complex filtering, which prunes nodes and edges that do not belong to three-dimensional simplicial complexes. We employ the same train/test split as Jo et al. (2022) for a fair comparison with baselines. Consistent with You et al. (2018), we employ the maximum mean discrepancy (MMD) to quantify the distribution differences across key graph statistics, including degree (Deg.), clustering coefficient (Clus.), and the number of occurrences of orbits with 4 nodes (Orbit). To provide a holistic evaluation, we further calculate the mean MMD across these metrics, which is reported in the Avg. column as the overall assessment index. A lower MMD signifies a closer alignment between the generated and evaluation datasets, indicating superior generative performance.
Baselines
We compare our model with prominent auto-regressive and one-shot graph generation approaches. Auto-regressive models sequentially build graphs by adding nodes and edges step-by-step. Benchmarks in this category include DeepGMG (Li et al., 2018), GraphRNN (You et al., 2018), GraphAF (Shi et al., 2020), and GraphDF (Luo et al., 2021). Conversely, one-shot models generate all nodes and edges simultaneously. We include GraphVAE (Simonovsky & Komodakis, 2018), GNF (Liu et al., 2019), GPrinFlowNet (Mo et al., 2024), EDP-GNN (Niu et al., 2020), GDSS (Jo et al., 2022), and DiGress (Vignac et al., 2023).
Sampling Quality
The results in Tab. 2 verify that HOG-Diff is not only suitable for molecular generation but also proficient in generic graph generation. These experiments demonstrate HOG-Diff’s ability to model the intricate interdependencies between nodes and edges effectively.
@@ 4.3 Ablations: Topological Guide Analysis
Noise 0.002 1.665 96.78
ZINC250k Peripheral 0.002 1.541 97.93
Cell 0.001 1.533 98.56
Figure 4: (Left) Training curves of the score-matching process. The entire process of HOG-Diff is divided into two stages, i.e., $K=2$, referred to as coarse and fine, respectively. The combined loss of these two stages is labelled as Coarse+Fine. (Right) Sampling results of various topological guides.
During the experiments, we observe that HOG-Diff exhibits superior performance on complex datasets such as QM9 and Zinc250k, but comparatively modest results on the Ego dataset. Visualizations and statistics in Apps. F and G indicate that Ego contains the fewest higher-order structures among the datasets analyzed, suggesting that the choice of guide plays a pivotal role in the effectiveness of generation. To validate this hypothesis, we conduct further ablations using different types of topological information as guides.
Specifically, we employ three types of guiders: structures derived from 2-cell filtering (Cell), peripheral structures obtained by removing cell components (Peripheral), and Gaussian random noise (Noise). Employing noise as the guide aligns with classical generation models which generate samples by denoising noisy data. Fig. 4 (plot) visualizes how the spectrum loss changes during the training process and demonstrates that training the proposed model converges faster compared to the classical method, which aligns with the theoretical results in Prop. 3.3.
The sampling results in Fig. 4 (table) show that both peripheral and noise guides are inferior compared to using cell structures as the guide, empirically supporting Prop. 3.4. This highlights that certain topological structures, such as cells, are more effective in guiding the generation, likely due to their higher-order connectivity and structural significance. These observations stress the importance of identifying and leveraging proper topological structures as guides, which play a critical role in steering the generative process toward meaningful outputs. Moreover, this property opens up promising avenues for exploring the guide’s potential as a tool to diagnose whether a specific component is integral and essential for the architecture. Furthermore, by systematically analyzing the impact of various guides, we can deepen our understanding of the interplay between structural characteristics and generative performance, thereby advancing the design of more effective graph generative models.
@@ 5 Related Works
We review graph generation methods along with higher-order generation. App. C presents a more detailed review.
@@
Graph generative models make great progress by exploiting the capacity of deep neural networks. These models typically generate nodes and edges either in an autoregressive manner or simultaneously, utilizing techniques such as variational autoencoders (VAE) (Jin et al., 2018; Simonovsky & Komodakis, 2018), recurrent neural networks (RNN) (You et al., 2018), normalizing flows (Zang & Wang, 2020; Shi et al., 2020; Luo et al., 2021), and generative adversarial networks (GAN) (De Cao & Kipf, 2018; Martinkus et al., 2022).
@@
A breakthrough in graph generative models has been marked by the recent progress in diffusion-based generative models (Niu et al., 2020). Recent models employ various strategies to enhance the generation of complex graphs, including capturing node-edge dependency (Jo et al., 2022), addressing discretization challenges (Vignac et al., 2023; Huang et al., 2023), exploiting low-to-high frequency generation curriculum (Mo et al., 2024), and improving computational efficiency through low-rank diffusion processes (Luo et al., 2023a). Recent studies have also enhanced diffusion-based generative models by incorporating diffusion bridge processes, i.e., processes conditioned on the endpoints (Wu et al., 2022; Boget et al., 2024; Jo et al., 2024). Despite these advances, existing methods either overlook or inadvertently disrupt higher-order structures during graph generation, or struggle to model the topological properties, as denoising the noisy samples does not explicitly preserve the intricate structural dependencies required for generating realistic graphs.
@@
Generative modelling uses higher-order information mostly in the form of hypergraphs. Models such as Hygene (Gailhard et al., 2024) and HypeBoy (Kim et al., 2024) aim to generate hypergraphs. Dymond (Zeno et al., 2021) focuses on higher-order motifs in dynamic graphs. To the best of our knowledge, we are the first to consider higher-order guides for graph generation.
@@ 6 Conclusion
We introduce HOG-Diff, a coarse-to-fine generation framework that explicitly exploits higher-order graph topology. It decomposes the complicated generation process into easier-to-learn sub-steps, which are implemented using a generalized OU bridge process. Our theoretical analysis justifies the effectiveness of HOG-Diff over classical diffusion approaches, which is validated by superior experimental results on both molecular and generic graph generation tasks. Our framework further promises to improve interpretability by enabling the analysis of different topological guides’ performance in the generation process. This work is a key step in topological diffusion models, highlighting the impact of higher-order features absent in data and opening ample room for future work.
Impact Statement

### v2

@@ 3.3 Theoretical Analysis
We now provide theoretical evidence for the efficacy of HOG-Diff, demonstrating that the proposed framework achieves faster convergence in score-matching and tighter reconstruction error bounds compared to standard graph diffusion. We experimentally verify our theories in Sec. 4.3.
Theorem 3 (Informal).
@@
Suppose the loss function $\ell^{(k)}(\bm{\theta})$ in Eq. 8 is $\beta$-smooth and satisfies the $\mu$-PL condition in the ball $B\left(\bm{\theta}_{0},R\right)$. Then, the expected loss at the $i$-th training iteration satisfies:
$\mathbb{E}\left[\ell^{(k)}(\bm{\theta}_{i})\right]\leq\left(1-\frac{b\mu^{2}}{\beta N(\beta N^{2}+\mu(b-1))}\right)^{i}\ell^{(k)}\left(\bm{\theta}_{0}\right),$ (9)
@@
Following Luo et al. (2023a), we define the expected reconstruction error at each generation process as $\mathcal{E}(t)=\mathbb{E}\left\|\bar{\bm{G}}_{t}-\widehat{\bm{G}}_{t}\right\|^{2}$, where $\bar{\bm{G}}_{t}$ represents the data reconstructed with the ground truth score $\nabla\log p_{t}(\cdot)$ and $\widehat{\bm{G}}_{t}$ denotes the data reconstructed with the learned score function $\bm{s}_{\bm{\theta}}$. Next, we establish that the reconstruction error in HOG-Diff is bounded more tightly than in classical graph generation models, thereby ensuring superior sample quality.
Theorem 4.
@@
where $\alpha(0)=C^{2}\ell^{(1)}(\bm{\theta})\int_{0}^{\tau_{1}}g_{1,s}^{4}\mathop{}\!\mathrm{d}{s}+C\mathcal{E}(\tau_{1})\int_{0}^{\tau_{1}}h_{1,s}^{2}\mathop{}\!\mathrm{d}{s}$, $\gamma(s)=C^{2}g_{1,s}^{4}\|\bm{s}_{\bm{\theta}}(\cdot,s)\|_{\mathrm{lip}}^{2}+C\|h_{1,s}\|_{\mathrm{lip}}^{2}$, and $h_{1,s}=\theta_{s}\left(1+\frac{2}{e^{2\bar{\theta}_{s:\tau_{1}}}-1}\right)$. Furthermore, we can derive that the reconstruction error bound of HOG-Diff is sharper than that of classical graph generation models.
The theorems above rely primarily on mild assumptions, such as smoothness and boundedness, without imposing strict conditions like the target distribution being log-concave or satisfying the log-Sobolev inequality. Their formal statements and detailed proofs are postponed to App. A.
@@ 4 Experiments
Figure 3: Visualization of molecular graphs at different stages of the reverse generative process. Model trained on the ZINC250k dataset.
We assess HOG-Diff against state-of-the-art baselines in both molecular and generic graph generation. Ablation studies are further conducted to analyze the impact of different topological guides. Complexity analysis and experimental settings are deferred to Apps. D and E, while App. F presents further results, including diffusion domain analysis, large-scale SBM experiments, the rationale for filtering choices, variance statistics, and visualizations.
Table 1: Comparison of different methods based on molecular datasets. We report the mean of 3 different runs. The best results for the first three metrics are highlighted in bold.
Method QM9 ZINC250k
[…]
Val. w/o
[…]
Val. w/o
[…]
GraphAF 0.021 5.625 74.43 88.64 86.59 0.044 16.023 68.47 98.64 99.99
GraphDF 0.064 10.928 93.88 98.58 98.54 0.177 33.546 90.61 99.63 100.00
GraphArm 0.002 1.220 90.25 95.62 70.39 0.055 16.260 88.23 99.46 100.00
MiCaM 0.001 1.045 99.93 93.89 83.25 0.166 31.495 100.00 88.48 99.98
GraphEBM 0.030 6.143 8.22 97.90 97.01 0.212 35.471 5.29 98.79 100.00
SPECTRE 0.163 47.960 87.30 35.70 97.28 0.109 18.440 90.20 67.05 100.00
GSDM 0.003 2.650 99.90 - - 0.017 12.956 92.70 - -
EDP-GNN 0.005 2.680 47.52 99.25 86.58 0.049 16.737 82.97 99.79 100.00
GDSS 0.003 2.900 95.72 98.46 86.27 0.019 14.656 97.01 99.64 100.00
DiGress 0.0005 0.360 99.00 96.66 33.40 0.082 23.060 91.02 81.23 100.00
MoFlow 0.017 4.467 91.36 98.65 94.72 0.046 20.931 63.11 99.99 100.00
CatFlow - 0.441 99.81 99.95 - - 13.211 99.95 99.99 -
DeFoG 0.0005 0.268 99.26 96.61 72.57 0.002 2.030 94.97 99.98 100.00
HOG-Diff 0.0003 0.172 98.74 97.10 75.12 0.001 1.633 98.56 99.96 99.53
@@ 4.1 Molecule Generation
Molecular design is a prominent application of graph generation. We conduct evaluations on two well-known molecular datasets: QM9 (Ramakrishnan et al., 2014) and ZINC250k (Irwin et al., 2012). Intermediate higher-order skeletons are extracted using 2-cell complex filtering; the rationale for this choice is discussed in Sec. F.3. We evaluate the quality of 10,000 generated molecules with five standard metrics as in Jo et al. (2022): Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) MMD (Costa & Grave, 2010), Fréchet ChemNet Distance (FCD) (Preuer et al., 2018), Validity without correction (Val. w/o corr.), Uniqueness (Uni.), and Novelty (Nov.) (Jo et al., 2022).
For benchmarking, we include various representative molecular generation models. Autoregressive models include GraphAF (Shi et al., 2020), GraphDF (Luo et al., 2021), and GraphArm (Kong et al., 2023); the fragment-based MiCaM (Geng et al., 2023) creatively leverage motif information. In contrast, the remaining methods adopt a one-shot generation paradigm: GraphEBM serves as an energy-based model, SPECTRE (Martinkus et al., 2022) and GSDM (Luo et al., 2023a) incorporate spectral conditioning within GAN and diffusion frameworks, respectively, while EDP-GNN (Niu et al., 2020), GDSS (Jo et al., 2022), and DiGress (Vignac et al., 2023) represent diffusion-based graph generation models. We also compare against advanced flow-based models, including MoFlow (Zang & Wang, 2020), CatFlow (Eijkelboom et al., 2024), and DeFoG (Qin et al., 2025).
We visualize the molecule generation process in Fig. 3 with more examples deferred to Sec. F.5. It can be observed that our model explicitly preserves higher-order structures during the generation process. Tab. 1 indicates that HOG-Diff consistently outperforms both auto-regressive and one-shot models. Notably, the dramatic decrease in NSPDK and FCD implies that HOG-Diff is able to generate molecules with data distributions close to those of real molecules in both chemical and graph spaces. Additional results on validity and variance are provided in Sec. F.4.
@@ 4.2 Generic Graph Generation
To display the topology distribution learning ability, we assess HOG-Diff over four common generic graph datasets: Community-small, Ego-small, Enzymes, and a larger-scale stochastic block model (SBM) dataset. The SBM dataset is presented separately in Sec. F.2 due to its distinct evaluation protocol. Intermediate higher-order skeletons are obtained via 3-simplicial complex filtering. We employ the same train/test split as Jo et al. (2022) for a fair comparison. Maximum mean discrepancy (MMD) is used to quantify the distribution differences across key graph statistics, including degree (Deg.), clustering coefficient (Clus.), and 4-node orbit counts (Orbit). A lower MMD signifies a closer alignment between the generated and evaluation datasets, suggesting superior generative performance. We also report the average MMD across all metrics as an overall indicator.
Table 2: Generation performance on generic graph datasets. Best bold and second-best underlined. Hyphen (-) indicates missing results in the original paper.
Method Community-small Enzymes Ego-small
@@
Figure 4: Training curves of the score-matching process. The entire process of HOG-Diff is divided into two stages, i.e., $K=2$, referred to as coarse and fine, respectively. The combined loss of these two stages is labelled as Coarse+Fine.
We compare the following graph generative models: DeepGMG (Li et al., 2018) and GraphRNN (You et al., 2018) are autoregressive models, while GraphVAE (Simonovsky & Komodakis, 2018), GNF (Liu et al., 2019), and GPrinFlowNet (Mo et al., 2024) are one-shot models. GraphAF, GraphDF, EDP-GNN, SPECTRE, GDSS, and DiGress are previously explained. The results in Tab. 2 verify that HOG-Diff is not only suitable for molecular generation but also proficient in generic graph generation, demonstrating its ability to effectively capture the intricate topological interdependencies.
@@ 4.3 Ablations: Topological Guide Analysis
During the experiments, we observe that HOG-Diff exhibits superior performance on complex datasets such as QM9 and ZINC250k, but comparatively modest results on the Ego-small dataset. Statistics and visualizations in Apps. E and F.5 reveal that Ego contains the fewest higher-order structures among the datasets analyzed, suggesting that the choice of guide plays a pivotal role in the effectiveness of generation. To validate this hypothesis, we conduct further ablations using different types of topological information as guides.
Table 3: Sampling results of various topological guides.
[…]
Val. w/o
@@
ZINC
250k
Noise 0.002 2.665 96.78
Periph. 0.002 2.641 97.93
Cell 0.001 1.633 98.56
We first compare two types of guides: structures derived from 2-cell filtering (Cell) and Gaussian random noise (Noise). Employing noise as a guide aligns with classical diffusion paradigms that generate samples by progressively denoising noisy data. Fig. 4 visualizes how the spectrum loss changes during the training process. It shows that our framework (red curve) converges faster than the classical method (blue curve), which is consistent with our theoretical results in Theorem 3.
In the sampling procedure, we further evaluate peripheral structures (Periph.), obtained by removing cell components, as guides. As shown in Tab. 3, both peripheral and noise guides perform worse than cell-based guides, providing empirical support for the tighter reconstruction error bound established in Theorem 4. These results indicate that certain topologies, particularly cells, are more effective in guiding generation, highlighting the importance of selecting appropriate topological structures to steer toward meaningful outputs. Moreover, this finding suggests that guides could serve as diagnostic tools for assessing whether specific topologies are essential to the architecture. Systematic analysis of different guides promises to enhance interpretability, clarify how topologies affect generation, and inform the design of more effective graph generative models.
@@ 5 Related Works
We review graph and higher-order generation methods, with a more detailed discussion in App. B.
@@
Graph generative models have made great progress by exploiting the capacity of deep neural networks. These models typically generate nodes and edges either in an autoregressive manner or simultaneously, utilizing techniques such as variational autoencoders (VAE) (Jin et al., 2018; Simonovsky & Komodakis, 2018), recurrent neural networks (RNN) (You et al., 2018), normalizing flows (Zang & Wang, 2020; Shi et al., 2020; Luo et al., 2021), and generative adversarial networks (GAN) (Martinkus et al., 2022).
@@
A breakthrough in graph generative models has been marked by the recent progress in diffusion-based generative models (Niu et al., 2020). Recent models employ various strategies to enhance the generation of complex graphs, including capturing node-edge dependency (Jo et al., 2022), addressing discretization challenges (Vignac et al., 2023; Huang et al., 2023), exploiting low-to-high frequency generation curriculum (Mo et al., 2024), and improving computational efficiency through low-rank diffusion processes (Luo et al., 2023a). CatFlow (Eijkelboom et al., 2024) and DeFoG (Qin et al., 2025) adopt flow matching as an alternative to diffusion, achieving more efficient generation. Recent studies have also enhanced diffusion-based generative models by incorporating diffusion bridge processes, i.e., processes conditioned on the endpoints (Wu et al., 2022; Boget et al., 2024; Jo et al., 2024). Despite these advances, existing methods either overlook or inadvertently disrupt higher-order structures during graph generation, or struggle to model the topological properties, as denoising the noisy samples does not explicitly preserve the intricate structural dependencies required for generating realistic graphs.
@@
Generative modeling uses higher-order information mostly in the form of motifs and hypergraphs. MiCaM (Geng et al., 2023) synthesizes molecules by iteratively merging motifs. HypeBoy (Kim et al., 2024) learns hypergraph representation through hyperedge filling, while Hygene (Gailhard et al., 2025) reduces hypergraph generation to bipartite graphs. To the best of our knowledge, we are the first to consider higher-order guides for graph generation.
@@ 6 Conclusion
We introduce HOG-Diff, a coarse-to-fine generation framework that explicitly exploits higher-order graph topology. It decomposes the complicated graph generation process into easier-to-learn sub-steps, which are implemented using a GOU bridge process. Our theoretical analysis justifies the effectiveness of HOG-Diff over classical diffusion-based approaches, which is further validated by superior experimental results on both molecular and generic graph generation tasks. This work is a key step in topological diffusion models, highlighting the impact of higher-order features absent in data and opening ample room for future work.
Ethics and Reproducibility Statement
@@
Experimental code related to this paper is provided in the Supplementary Material. Detailed theoretical derivations are provided in App. A. The complete architecture of HOG-Diff, along with the training objectives and sampling procedures, is described in App. C. Details of the datasets, preprocessing steps, and experimental settings are provided in App. E. Additional experimental results, ablation studies, and visualizations can be found in App. F.
In this work, Large Language Models (LLMs) are used solely for language polishing.
Acknowledgments
T. Birdal acknowledges support from the Engineering and Physical Sciences Research Council [grant EP/X011364/1]. T. Birdal was supported by a UKRI Future Leaders Fellowship [grant number MR/Y018818/1] as well as a Royal Society Research Grant RG/R1/241402.